# Brownian dynamics simulations of electric double-layer capacitors with tunable metallicity

### v1

@@ II.3.3 Simulation details
The coupled equations Eq. 38 are solved numerically using the MetalWalls simulation package marin-lafleche_metalwalls_2020; coretti_metalwalls_2022, in which we implemented the Euler–Maruyama integrator, a rescaling of the 2D Ewald summation by the permittivity $\varepsilon_{\rm s}$ to compute $U_{\rm Coul}^{\rm ions}$ (see Eq. 32) and the new term $\delta U_{\rm Coul}^{\rm ions}$ (see Eq. 46) to account for the effect of the implicit Thomas-Fermi electrodes, as well as the corresponding forces.
The parameters for the short-range interactions (see Section II.3.1) are identical for cations, anions and wall atoms, namely $\sigma_{ij}=\sigma_{\mathrm{w}}=5$ Å and $\epsilon_{ij}=\epsilon_{\mathrm{w}}=2.477$ kJ/mol. The cut-off for ion-ion interactions is thus $r^{*}=5.61$ Å, while that for the Steele potential is $z^{*}=4.93$ Å. For the definition of the latter, the graphite structure corresponds to $\rho_{\rm surf}=0.38$ Å-2 and $\Delta=3.354$ Å. The repulsive walls (explicit atom sites or zero of the Steele potential) are placed at the same position as the dielectric interface, i.e. $z=\pm L/2$.
@@
For BD simulations, we consider ions with a diffusion coefficient $D_{\pm}=1.12\times 10^{-9}$ m2s-1 and a temperature $T=298$ K. The overdamped Langevin equations 38 are integrated with a time step of 5 fs. The systems are first equilibrated for 10 ns, before production runs of 40 ns, with electrode charges (computed from Eq. II.2.3) sampled every 0.25 ps and ion positions every 0.5 ps. The reported uncertainty on ionic concentration profiles is the standard error over 10 consecutive blocks of the trajectory. The differential capacitance is computed from the variance of the charge distribution (first line in Eq. 37). The corresponding uncertainty is obtained as in Ref. scalfi_charge_2020 following Zwanzig and Ailawadi zwanzig_statiscal_1969: The standard error estimated as $\langle\delta Q^{2}\rangle\times\sqrt{4\tau_{\rm c}/\tau_{\rm s}}$ where $\tau_{\rm s}$ is the sampling time and $\tau_{\rm c}$ is the correlation time computed as $\tau_{\rm c}=\int_{0}^{\infty}\langle\delta Q(t)\delta Q(0)\rangle^{2}/\langle\delta Q^{2}\rangle^{2}\,\mathrm{d}t$ . The uncertainty over the integral capacity is computed via the same method with $\langle Q\rangle\times\sqrt{4\tau^{\prime}_{\rm c}/\tau_{\rm s}}$ with $\tau^{\prime}_{\rm c}=\int_{0}^{\infty}\langle\delta Q(t)\delta Q(0)\rangle/\langle\delta Q^{2}\rangle\,\mathrm{d}t$. We refer the reader to Ref. pireddu_impedance_2024 for more details on the computation of these correlation times.
@@ III Results
We now demonstrate the relevance of the proposed implicit electrode description, by first validating its predictions for the force on the ions as a function of the ion-surface distance in simple geometries: we compare our results with (semi-)analytical for isolated ions near electrodes for various $l_{\rm TF}$ in Section III.0.1) and with numerical results under periodic boundary conditions for ion pairs next to perfect metals ($l_{\rm TF}=0$) in Section III.0.2. We then turn in Section III.0.3 to the equilibrium ionic density profiles and compare our results with Brownian Dynamics simulations with an explicit electrode model for perfect metals proposed in Ref. cats_capacitance_2022, before using our model that also applies to finite $l_{\rm TF}$ to discuss the effect of the Thomas-Fermi screening length on the ionic density profiles in Section III.0.4. Section III.0.5 finally discusses the capacitance as a function of $l_{\rm TF}$ and compares, when possible, several levels of description.
@@ III.0.1 Isolated ions
where $d_{i}$ is the distance of ion $i$ from the nearest electrode. Eq. 49 was first derived by Kornyshev et al. (see Eqs. 6 and following of Ref. kornyshev_image_1977) and used as reference in several subsequent works (see in e.g. Eq. S19 of Ref. schlaich_electronic_2022 or more recently Eq. 69 of Ref. hedley_what_2025). This superposition approximation assumes in particular that $d_{i}\ll L$ for all ions and that the lateral dimensions are sufficiently large $d_{i}\ll L_{x},L_{y}$. Eq. 49 reduces to the well-known limits:
@@ III.0.2 Ion pair next to a perfect metal
We then compare the present method with the existing one for perfect metals ($l_{\rm TF}=0$) described in Ref. telles_efficient_2024 for a simple system where an ion pair is approached from the same electrode. Specifically, we consider a pair with charges +e+e at $(0,0,-\frac{L}{2}+d)$ and e-e at $(x,0,-\frac{L}{2}+d)$ with $x=0.714$ Å, in vacuum. The box dimensions are $L_{x}=67.69$ Å, $L_{y}=36.64$ Å, $L=39.72$ Å. Fig. 3 reports the components of the electrostatic force on the positive charge in the directions perpendicular and parallel to the electrodes as a function of the distance $d$ from them (the force on the negative charge is simply equal for the $z$ component and opposite for the $x$ one). The results for both components as a function of distance from the interface are in excellent agreement with the ones obtained by the method of Ref. telles_efficient_2024. This proves that the present approach, which is implemented as a correction to the case of ions in the absence of walls under 2D periodic boundary conditions (see Eq. 44) correctly describes the effect of the interface in the case of a perfect metal ($l_{\rm TF}=0$).
Figure 3: Electrostatic force in the directions perpendicular (a) and parallel (b) to the implicit electrodes as a function of the distance $d$ from them, for a pair of ions with charges +e+e at $(0,0,-\frac{L}{2}+d)$ and e-e at $(x,0,-\frac{L}{2}+d)$ with $x=0.714$ Å, in vacuum, for perfect metals ($l_{\rm TF}=0$). The reported force is that on the positive charge; the force on the negative charge is simply equal for the $z$ component and opposite for the $x$ one. The numerical results with the present method (lines) are compared with that obtained with the method of Ref. telles_efficient_2024 (symbols), which only applies to perfect metals.
@@ III.0.3 Comparison with explicit electrodes and implicit solvent
After the above validation for specific configurations of the ions by comparison with previous analytical and numerical results, we now turn to the study of equilibrium properties of capacitors obtained by Brownian dynamics simulations. We compare the results obtained with the present model describing both the polar solvent and the electrodes implicitly via the Green’s functions to that obtained in the case of perfect metals ($l_{\rm TF}=0$) by Cats et al. using explicit electrodes cats_capacitance_2022. In this work, the authors computed the forces on ions to be used in Brownian dynamics (in their case underdamped Langevin dynamics) by adapting the method introduced for molecular dynamics simulations with explicit electrode atoms maintained at a constant potential (with a potential difference between the two electrodes) and an explicit electrolyte (ions and solvent molecules). They suggested that for perfect metals, an implicit solvent can be modeled by simultaneously rescaling the ionic charges by a factor $1/\sqrt{\varepsilon_{\rm s}}$ and the voltage by a factor $\sqrt{\varepsilon_{\rm s}}$. While this rescaling in fact amounts to considering that the electrode atoms are also embedded in a medium with permittivity $\varepsilon_{\rm s}$, this approach resulted in an excellent agreement for the considered systems (monovalent ions in a very polar solvent) and voltages with classical Density Functional Theory calculations.
Here, we consider exactly the same system as in Ref. cats_capacitance_2022, namely a capacitor consisting of two graphite electrodes separated by $L=39.72$ Å. Each electrode is modeled by a single atomic layer of graphite with unit cell parameter $a=2.46$ Å and total dimensions $L_{x}\times L_{y}=67.69\times 36.64$ Å2, with a total of 960 atoms. We compare the results of overdamped Langevin dynamics simulations (Eq. 38) with forces computed as in Ref. cats_capacitance_2022 or with the present method, which does not rely on explicit electrode atoms to compute the electrostatic forces on the ions. For the short-range interactions, we consider either explicit atomic sites (interacting with the ions via pairwise WCA potentials, see Eq. 39) to isolate the effect of the electrostatic interactions, or implicit walls (see Eqs. 41-43), to isolate the effect of the short-range ones.
Figure 4: Equilibrium anion (red solid lines) and cation (blue dashed lines) density profiles from Brownian dynamics simulations of ions in an implicit solvent between perfect metals ($l_{\rm TF}=0$), with (a) the explicit electrode model from Ref. cats_capacitance_2022 (see text), (b) the present model of implicit electrodes and short-range interactions computed with the same atomic sites as in the previous panel and (c) the present fully implicit model where short-range interactions are computed via the Steele potential (see Eqs. 41-43). In all panels, the density profiles are shown for $\Delta\Psi=0$, $0.1$ and $0.2$ V from light to dark lines, respectively.
@@
The agreement between the various methods show that the same prediction for the ionic density profiles can be obtained with a significantly simpler model and, accordingly, a reduced computational cost. Indeed, for the considered system with a single layer of electrode atoms in each electrode (as in Ref. cats_capacitance_2022), with a total of 1920 electrode atoms in the system, the computational cost of electrostatic interactions is reduced by a factor of $\sim 6$ (all comparisons reported here are obtained from simulations with a single CPU on the same computer). For a system with three atomic planes in each electrodes, the speed-up is by a factor of $\sim 60$. For the same systems, the speed-up achieved using an implicit wall for short-range interactions instead of explicit sites is of approximately 25 and 80, respectively. For these systems, the computation of electrostatic interactions represents $\sim 95$% of the cost per step in the fully implicit model. Finally, while we compared the present implicit electrode model to one with explicit electrodes, in both cases the solvent is treated implicitly.
The speed-up with respect to a fully atomistic simulation is even more dramatic. As an example, for a system considered in Ref. pireddu_impedance_2024 with a 0.5 M NaCl aqueous solution between model gold electrodes separated by a distance $L\approx 5$ nm (with 40 ions, 2160 water molecules and 3240 electrode atoms), the present BD simulations (with only 40 ions and implicit solvent and electrodes) is $\approx 6$ times faster on a single CPU than the MD simulations on a GPU. The present model only approximately describes the effect of the solvent, neglecting in particular subtle effects that may occur in the first molecular layers at the interface, and a detailed comparison remains out of the scope of the present work (see some discussion in Section IV). Nevertheless, it offers a computationally efficient approach to investigate the effect of the TF screening length on the properties of capacitors with dimensions much larger than the ones currently within reach of MD simulations. This includes the possibility to consider lower salt concentrations (typically above 0.1 mol/L in MD simulations), by extending the lateral dimensions to ensure a sufficient number of ions in the system, or larger inter-electrode distances (typically limited to 10-20 nm in MD simulations). Importantly, the reduced computational cost per step and the larger time steps used in Brownian dynamics offer the possibility to also investigate the dynamics over much longer times scales (typically limited to a few tens of ns in MD simulations).
@@ III.0.4 Effect of the Thomas-Fermi screening length
We now use the present implicit model to investigate the effect of the Thomas-Fermi screening length on the ionic distributions and the capacitance. We recall that this was not possible with previous methods based on Green’s functions, which only applied so far to the perfect metal case girotto_simulations_2017; dos_santos_simulations_2017; telles_efficient_2024. In the following, we consider the same system as in Section III.0.3, except for the variable $l_{\rm TF}$.
@@ IV Conclusion and perspectives
The present model is limited to parallel plate capacitors, but can be used to investigate the effect of the TF screening length on the properties of capacitors with dimensions much larger than the ones currently within reach of molecular dynamics simulations. This would offer the possibility to consider lower salt concentrations (typically above 0.1 mol/L in MD simulations), by extending the lateral dimensions to ensure a sufficient number of ions in the system, or larger inter-electrode distances (typically limited to 10-20 nm in MD simulations). Importantly, the reduced computational cost per step and the larger time steps used in Brownian dynamics offer the possibility to also investigate the dynamics over much longer times scales (typically limited to a few tens of ns in MD simulations). It would in particular be interesting to derive the fluctuation-dissipation relation for the frequency-dependent admittance (inverse of the impedance), already successfully applied in MD simulations to link the electrochemical response with the dynamics within the liquid slab pireddu_frequency-dependent_2023; pireddu_impedance_2024, in the case of Brownian dynamics, for example following the work of Ref. hoang_ngoc_minh_frequency_2023 for the frequency-dependent conductivity of confined electrolytes. Beyond the study of these effects, the present BD approach can provide reference data to assess the predictions of simpler models such as PNP theory and its extensions, e.g. the ones mentioned in the introduction.
@@
Future possible directions also pertain to the description of the solvent. The effective interaction between an ion embedded in a dielectric and a Thomas-Fermi metal (see Fig. 2b) only partly reflects the actual potential of mean force (PMF), which could in principle be computed from all atoms MD simulations with constant-potential TF electrodes scalfi_semiclassical_2020. While a comparison of the present model with an explicit solvent is beyond the scope of the present work, we might anticipate that features such as the molecular layering of the solvent can lead to differences with the predictions of the present implicit model. Some of them can be captured in an effective way by tuning the position of the interface between the metal and the implicit solvent, using the concept of Dielectric Dividing Surface introduced by Netz an co-workers schlaich_water_2016 (see also Ref. cox_dielectric_2022), as illustrated for the frequency-dependent permittivity of nanocapacitors pireddu_frequency-dependent_2023; pireddu_impedance_2024.
@@
However others, in particular related to the solvation of interfacial ions, might require more elaborate descriptions such as molecular Density Functional Theory jeanmairet_molecular_2013; jeanmairet_study_2019; nair_ions_2025. Such molecular descriptions would however not allow to simply express the many-body effective potential between ions. This could however be achieved by introducing the solvent polarization as an additional field in the free energy functional and treating this polarization at the Born-Oppenheimer level, as the electron charge density. In fact, the present model can be recovered in such a perspective using the simplest functional corresponding to a linear dielectric medium characterized by its permittivity. More elaborate functionals for the solvent (see e.g. Refs. berthoumieux_dielectric_2019; becker_dielectric_2025; hedley_what_2025) could then be introduced naturally at this level of description. In practice, a simpler approach would be to introduce molecular features through a one-body ion-wall PMF, derived or inspired from MD simulations loche_effects_2022. Such terms (e.g. with damped oscillations to capture the solvent layering), have already been used in extensions of Poisson-Boltzmann theory horinek_specific_2007 or in simulations of electrokinetic phenomena joly_liquid_2006 and could straightforwardly be included in the present BD simulations via additional contributions to Eq. 41.

### v2

@@ II.3.3 Simulation details
The coupled equations Eq. 38 are solved numerically using the MetalWalls simulation package Marin-Laflèche et al. (2020); Coretti et al. (2022), in which we implemented the Euler–Maruyama integrator, a rescaling of the 2D Ewald summation by the permittivity $\varepsilon_{\rm s}$ to compute $U_{\rm Coul}^{\rm ions}$ (see Eq. 32) and the new term $\delta U_{\rm Coul}^{\rm ions}$ (see Eq. 46) to account for the effect of the implicit Thomas-Fermi electrodes, as well as the corresponding forces.
The parameters for the short-range interactions (see Section II.3.1) are identical for cations, anions and wall atoms, namely $\sigma_{ij}=\sigma_{\mathrm{w}}=5$ Å and $\epsilon_{ij}=\epsilon_{\mathrm{w}}=2.477$ kJ/mol. The cut-off for ion-ion interactions is thus $r^{*}=5.61$ Å, while that for the Steele potential is $z^{*}=4.92$ Å. For the definition of the latter, the graphite structure corresponds to $\rho_{\rm surf}=0.38$ Å-2 and $\Delta=3.354$ Å. The repulsive walls (explicit atom sites or zero of the Steele potential) are placed at the same position as the dielectric interface, i.e. $z=\pm L/2$.
@@
For BD simulations, we consider ions with a diffusion coefficient $D_{\pm}=1.12\times 10^{-9}$ m2s-1 and a temperature $T=298$ K. The overdamped Langevin equations 38 are integrated with a time step of 5 fs. The systems are first equilibrated for 100 ns, before production runs of 400 ns, with electrode charges (computed from Eq. II.2.3) sampled every 0.25 ps and ion positions every 0.5 ps. The reported uncertainty on ionic concentration profiles is the standard error over 10 consecutive blocks of the trajectory. The differential capacitance is computed from the variance of the charge distribution (first line in Eq. 37). The corresponding uncertainty is obtained as in Ref. 84 following Zwanzig and Ailawadi Zwanzig and Ailawadi (1969): The standard error estimated as $\langle\delta Q^{2}\rangle\times\sqrt{4\tau_{\rm c}/\tau_{\rm s}}$ where $\tau_{\rm s}$ is the sampling time and $\tau_{\rm c}$ is the correlation time computed as $\tau_{\rm c}=\int_{0}^{\infty}\langle\delta Q(t)\delta Q(0)\rangle^{2}/\langle\delta Q^{2}\rangle^{2}\,\mathrm{d}t$ . The uncertainty over the integral capacity is computed via the same method with $\langle Q\rangle\times\sqrt{4\tau^{\prime}_{\rm c}/\tau_{\rm s}}$ with $\tau^{\prime}_{\rm c}=\int_{0}^{\infty}\langle\delta Q(t)\delta Q(0)\rangle/\langle\delta Q^{2}\rangle\,\mathrm{d}t$. We refer the reader to Ref. 77 for more details on the computation of these correlation times.
@@ III Results
We now demonstrate the relevance of the proposed implicit electrode description, by first validating its predictions for the force on the ions as a function of the ion-surface distance in simple geometries: we compare our results with (semi-)analytical for isolated ions near electrodes for various $l_{\rm TF}$ in Section III.0.1) and with numerical results under periodic boundary conditions for ion pairs next to perfect metals ($l_{\rm TF}=0$) in Section III.0.2. We then turn in Section III.0.3 to the equilibrium ionic density profiles and compare our results with Brownian Dynamics simulations with an explicit electrode model for perfect metals proposed in Ref. 17, before using our model that also applies to finite $l_{\rm TF}$ to discuss the effect of the Thomas-Fermi screening length on the ionic density profiles in Section III.0.4. Section III.0.5 finally discusses the capacitance as a function of $l_{\rm TF}$ and compares, when possible, several levels of description.
@@ III.0.1 Isolated ions
where $d_{i}$ is the distance of ion $i$ from the nearest electrode. Eq. 49 was first derived by Kornyshev et al. (see Eqs. 6 and following of Ref. 51) and used as reference in several subsequent works (see in e.g. Eq. S19 of Ref. 86 or more recently Eq. 69 of Ref. 34). This superposition approximation assumes in particular that $d_{i}\ll L$ for all ions and that the lateral dimensions are sufficiently large $d_{i}\ll L_{x},L_{y}$. Eq. 49 reduces to the well-known limits:
@@ III.0.2 Ion pair next to a perfect metal
We then compare the present method with the existing one for perfect metals ($l_{\rm TF}=0$) described in Ref. 93 for a simple system where an ion pair is approached from the same electrode. Specifically, we consider a pair with charges +e+e at $(0,0,-\frac{L}{2}+d)$ and e-e at $(x,0,-\frac{L}{2}+d)$ with $x=0.714$ Å, in vacuum. The box dimensions are $L_{x}=67.69$ Å, $L_{y}=36.64$ Å, $L=39.72$ Å. Fig. 3 reports the components of the electrostatic force on the positive charge in the directions perpendicular and parallel to the electrodes as a function of the distance $d$ from them (the force on the negative charge is simply equal for the $z$ component and opposite for the $x$ one). The results for both components as a function of distance from the interface are in excellent agreement with the ones obtained by the method of Ref. 93. This proves that the present approach, which is implemented as a correction to the case of ions in the absence of walls under 2D periodic boundary conditions (see Eq. 44) correctly describes the effect of the interface in the case of a perfect metal ($l_{\rm TF}=0$).
Figure 3: Electrostatic force in the directions perpendicular (a) and parallel (b) to the implicit electrodes as a function of the distance $d$ from them, for a pair of ions with charges +e+e at $(0,0,-\frac{L}{2}+d)$ and e-e at $(x,0,-\frac{L}{2}+d)$ with $x=0.714$ Å, in vacuum, for perfect metals ($l_{\rm TF}=0$). The reported force is that on the positive charge; the force on the negative charge is simply equal for the $z$ component and opposite for the $x$ one. The numerical results with the present method (lines) are compared with that obtained with the method of Ref. 93 (symbols), which only applies to perfect metals.
@@ III.0.3 Comparison with explicit electrodes and implicit solvent
After the above validation for specific configurations of the ions by comparison with previous analytical and numerical results, we now turn to the study of equilibrium properties of capacitors obtained by Brownian dynamics simulations. We compare the results obtained with the present model describing both the polar solvent and the electrodes implicitly via the Green’s functions to that obtained in the case of perfect metals ($l_{\rm TF}=0$) by Cats et al. using explicit electrodes 17. In this work, the authors computed the forces on ions to be used in Brownian dynamics (in their case underdamped Langevin dynamics) by adapting the method introduced for molecular dynamics simulations with explicit electrode atoms maintained at a constant potential (with a potential difference between the two electrodes) and an explicit electrolyte (ions and solvent molecules). They suggested that for perfect metals, an implicit solvent can be modeled by simultaneously rescaling the ionic charges by a factor $1/\sqrt{\varepsilon_{\rm s}}$ and the voltage by a factor $\sqrt{\varepsilon_{\rm s}}$. While this rescaling in fact amounts to considering that the electrode atoms are also embedded in a medium with permittivity $\varepsilon_{\rm s}$, this approach resulted in an excellent agreement for the considered systems (monovalent ions in a very polar solvent) and voltages with classical Density Functional Theory calculations.
Here, we consider exactly the same system as in Ref. 17, namely a capacitor consisting of two graphite electrodes separated by $L=39.72$ Å. Each electrode is modeled by a single atomic layer of graphite with unit cell parameter $a=2.46$ Å and total dimensions $L_{x}\times L_{y}=67.69\times 36.64$ Å2, with a total of 960 atoms. We compare the results of overdamped Langevin dynamics simulations (Eq. 38) with forces computed as in Ref. 17 or with the present method, which does not rely on explicit electrode atoms to compute the electrostatic forces on the ions. For the short-range interactions, we consider either explicit atomic sites (interacting with the ions via pairwise WCA potentials, see Eq. 39) to isolate the effect of the electrostatic interactions, or implicit walls (see Eqs. 41-43), to isolate the effect of the short-range ones.
Figure 4: Equilibrium anion (red solid lines) and cation (blue dashed lines) density profiles from Brownian dynamics simulations of ions in an implicit solvent between perfect metals ($l_{\rm TF}=0$), with (a) the explicit electrode model from Ref. 17 (see text), (b) the present model of implicit electrodes and short-range interactions computed with the same atomic sites as in the previous panel and (c) the present fully implicit model where short-range interactions are computed via the Steele potential (see Eqs. 41-43). In all panels, the density profiles are shown for $\Delta\Psi=0$, $0.1$ and $0.2$ V from light to dark lines, respectively.
@@
The agreement between the various methods show that the same prediction for the ionic density profiles can be obtained with a significantly simpler model and, accordingly, a reduced computational cost. Indeed, for the considered system with a single layer of electrode atoms in each electrode (as in Ref. 17), with a total of 1920 electrode atoms in the system, the computational cost of electrostatic interactions is reduced by a factor of $\sim 6$ (all comparisons reported here are obtained from simulations with a single CPU on the same computer). For a system with three atomic planes in each electrodes, the speed-up is by a factor of $\sim 60$. For the same systems, the speed-up achieved using an implicit wall for short-range interactions instead of explicit sites is of approximately 25 and 80, respectively. For these systems, the computation of electrostatic interactions represents $\sim 95$% of the cost per step in the fully implicit model. Finally, while we compared the present implicit electrode model to one with explicit electrodes, in both cases the solvent is treated implicitly.
The speed-up with respect to a fully atomistic simulation is even more dramatic. As an example, for a system considered in Ref. 77 with a 0.5 M NaCl aqueous solution between model gold electrodes separated by a distance $L\approx 5$ nm (with 40 ions, 2160 water molecules and 3240 electrode atoms), the present BD simulations (with only 40 ions and implicit solvent and electrodes) is $\approx 6$ times faster on a single CPU than the MD simulations on a GPU. The present model only approximately describes the effect of the solvent, neglecting in particular subtle effects that may occur in the first molecular layers at the interface, and a detailed comparison remains out of the scope of the present work (see some discussion in Section IV). Nevertheless, it offers a computationally efficient approach to investigate the effect of the TF screening length on the properties of capacitors with dimensions much larger than the ones currently within reach of MD simulations. This includes the possibility to consider lower salt concentrations (typically above 0.1 mol/L in MD simulations), by extending the lateral dimensions to ensure a sufficient number of ions in the system, or larger inter-electrode distances (typically limited to 10-20 nm in MD simulations). Importantly, the reduced computational cost per step and the larger time steps used in Brownian dynamics offer the possibility to also investigate the dynamics over much longer times scales (typically limited to a few tens of ns in MD simulations).
@@ III.0.4 Effect of the Thomas-Fermi screening length
We now use the present implicit model to investigate the effect of the Thomas-Fermi screening length on the ionic distributions and the capacitance. We recall that this was not possible with previous methods based on Green’s functions, which only applied so far to the perfect metal case Girotto et al. (2017); Dos Santos et al. (2017); Telles et al. (2024). In the following, we consider the same system as in Section III.0.3, except for the variable $l_{\rm TF}$.
@@ IV Conclusion and perspectives
The present model is limited to parallel plate capacitors, although the various steps of Section II could in principle be extended to undulating surfaces, following Ref. 79 where Green’s functions were derived for undulating dielectric membranes. But it can be used to investigate the effect of the TF screening length on the properties of capacitors with dimensions much larger than the ones currently within reach of molecular dynamics simulations. This would offer the possibility to consider lower salt concentrations (typically above 0.1 mol/L in MD simulations), by extending the lateral dimensions to ensure a sufficient number of ions in the system, or larger inter-electrode distances (typically limited to 10-20 nm in MD simulations). Importantly, the reduced computational cost per step and the larger time steps used in Brownian dynamics offer the possibility to also investigate the dynamics over much longer times scales (typically limited to a few tens of ns in MD simulations). It would in particular be interesting to derive the fluctuation-dissipation relation for the frequency-dependent admittance (inverse of the impedance), already successfully applied in MD simulations to link the electrochemical response with the dynamics within the liquid slab Pireddu and Rotenberg (2023); Pireddu et al. (2024), in the case of Brownian dynamics, for example following the work of Ref. 35 for the frequency-dependent conductivity of confined electrolytes. Beyond the study of these effects, the present BD approach can provide reference data to assess the predictions of simpler models such as PNP theory and its extensions, e.g. the ones mentioned in the introduction.
In the present work, we only considered systems without redox reactions at the electrode surfaces. In this purely capacitive case, one can assume that the electrolyte is overall neutral (hence that the two electrode charges are opposite) to obtain simple expressions for the electrode charge (Eq. II.2.3) and the many-body effective potential (Eq. 31). However many steps of the derivation in Section II do not rely on this assumption. In particular, Eq. 30 remains valid and one could obtain an expression when the electrolyte is not neutral and its charge is compensated by the net charge of the electrodes. The result could then be used in hybrid simulations where the Brownian dynamics of ions is combined with a Monte Carlo scheme for tentative instantaneous charge transfers between ions and electrodes, as done in Ref. 24 for perfect metals. Such an extension would allow to investigate the effect of the Thomas-Fermi length e.g. the relaxation of the electric double layer after an electron transfer Grün et al. (2004) or on the correlation between electrodes in electrochemical cells Huang et al. (2023).
Even though the TF model was successfully used to predict the effect of the screening length on a variety of interfacial properties, the accuracy of its predictions for a specific material can be limited, especially for properties that crucially depend on short-range features (expectedly so, since it is the low-wavevector and low-frequency limit of Lindhard theory of screening). Subtle effects, such as the electron spillover at the surface of the metal, which can enhance the capacitance (sometimes described as negative contributions to the inverse capacitance, e.g. in jellium models Schmickler and Henderson (1984); Leiva and Schmickler (1987)), could be approximately captured at the present level by a careful definition of an effective sharp interface between the jellium-like metal and its neighborhood (vacuum or liquid), see e.g. Refs. 84; 72. In order to go beyond, one could follow the steps described in Section II with a better kinetic energy functional. Adding the von Weizsäcker correction, that involves the gradient of the electronic density Weizsäcker (1935); Benguria et al. (1981), to the TF term Eq. 5 would be a natural step in this direction.
@@
Future possible improvements also pertain to the description of the solvent. The effective interaction between an ion embedded in a dielectric and a Thomas-Fermi metal (see Fig. 2b) only partly reflects the actual potential of mean force (PMF), which could in principle be computed from all atoms MD simulations with constant-potential TF electrodes Scalfi et al. (2020a). While a comparison of the present model with an explicit solvent is beyond the scope of the present work, we might anticipate that features such as the molecular layering of the solvent can lead to differences with the predictions of the present implicit model. Some of them can be captured in an effective way by tuning the position of the interface between the metal and the implicit solvent, using the concept of Dielectric Dividing Surface introduced by Netz an co-workers Schlaich et al. (2016) (see also Ref. 21), as illustrated for the frequency-dependent permittivity of nanocapacitors Pireddu and Rotenberg (2023); Pireddu et al. (2024).
@@
However others, in particular related to the solvation of interfacial ions, might require more elaborate descriptions such as molecular Density Functional Theory Jeanmairet et al. (2013, 2019); Nair et al. (2025a). Such molecular descriptions would however not allow to simply express the many-body effective potential between ions. This could however be achieved by introducing the solvent polarization as an additional field in the free energy functional and treating this polarization at the Born-Oppenheimer level, as the electron charge density. In fact, the present model can be recovered in such a perspective using the simplest functional corresponding to a linear dielectric medium characterized by its permittivity. More elaborate functionals for the solvent (see e.g. Refs. 10; 8; 34) could then be introduced naturally at this level of description. In practice, a simpler approach would be to introduce molecular features through a one-body ion-wall PMF, derived or inspired from MD simulations Loche et al. (2022). Such terms (e.g. with damped oscillations to capture the solvent layering), have already been used in extensions of Poisson-Boltzmann theory Horinek and Netz (2007) or in simulations of electrokinetic phenomena Joly et al. (2006) and could straightforwardly be included in the present BD simulations via additional contributions to Eq. 41.